• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 23:56:33 on February 2,2002

    Entry number 81593

    keyword=shift summary

    16:00 at 21.829 degrees, run 1991, beam unsteady
    16:25 1992, beam back

    EOR and BOR comments are incorrect for runs 1991 and 1992 (the angles were not updated).

    raster currents (for 5 pass at 5.7 GeV) are 13.41 amps on MRA1H00H and 23.35 amps on MRA1H00V. This gives a spot size 2 mm by 2 mm peak to peak.
    16:40 at 22.361 degrees, run 1993. beam unsteady
    17:15 MCC phoned about a H2 sensor (on CH004 of the FSD overview screens) which is faulted. We do not see a target problem and called in K. McCormick. MCC says in the past this HAD been masked but is now not. Jack Segal was paged and asked us to verify the GasMaster was not alarming (it is not). Dave Meekins is coming in to take a look and we have requested a hall access.
    17:30 MCC phoned. They believe the fault is in the dump. MCC has called in AES to take a look.
    17:55 MCC calls again. The sensor is running about 12% and the trip point should be 50% so they are going to mask the sensor and try to re-establish beam.
    18:50 beam is stable at 10 uAmps and we are starting run 1994 at 22.361 degrees.

    19:00 MCC cannot give more than 19 uAmps. MCC declined 30 uAmps, worried that it would trip off Hall B. We will run for less statistics.

    19:05 changing to 23.893 degrees, starting run 1995. DAQ crashed (ER1 disconnected). Bogdan does the kcoda and reboots daq.
    19:30 moved to 23.428 degrees. Started run 1997.
    19:50 changed to 23.958. Started run 1998.

    20:00 requested access to change to DVCS test plan.
    20:50 Q2 hrsL power supply tripped off. Go to window controlling power supply and find no fault conditions. Pushed reset button and typed in 3.5 GeV/c for momentum. Supply is now ramping back up.

    21:50 asked for beam back at 1 uAmp to start.

    24:00 MCC has had trickles of beam in the hall, but nothing for us to work with. MCC calls and asks if they can tune with 10 uA to setup the beam line. Eugene agrees.